# SADA: Semantic adversarial unsupervised domain adaptation for Temporal Action Localization

### v2

@@ B.2 CharadesEgo scenario
Third-to-egocentric setup: Based on this dataset, we are the first to present a benchmark on domain adaptation for Temporal Action Localization. Our presented setup, concretely, explores the adaptation of third-person videos to egocentric ones. This involves an extreme domain shift. For this, following the work of [12], we define the third-person videos as Source domain and the egocentric ones as Target domain. This setup presents an important limitation that we find is critical to make this benchmark viable in this complex task. Concretely, observe that class actions are mostly defined as NLP descriptions, which creates very concrete actions, making it cumbersome to adapt classes across domains. We argue that one of the main reasons is the lack of sufficient data for each of the actions. Consequently, we rely on the originally extended annotations from [13] and propose to use instead the verbs of each of the corresponding action labels. This is, for a given action holding some clothes, we consider this action with its verb hold. One final consideration is that this still results in underrepresented verb classes which we fix following a similar approach to that presented by [61] and consider only the 10 majority classes being take, put, sit, walk, hold, stand, open, drink, smile and laugh. This still results in a very challenging setup as observed in Table II of the main text, which indicates that none of the state-of-the-art methods tested attain more than mAP. Moreover, this setup retains of the original annotations.
Appendix C Unsupervised domain adaptation baselines
In this section, we describe in greater detail the different baselines that we use in Sec. IV.C of the main text, to establish a fair comparison with our work, SADA. Concretely, we first describe the adaptations to our architecture of multiple UDA methods like DANN [3], ADDA [40], WDGRL [42] or MSTN [2]. We then describe the adaptation of the closest related work to SADA, SSTDA [7] and for comparison purposes, the adaptation of one of the latest video action recognition methods, TranSVAE [65] to the task of Temporal Action Localization.
@@ Appendix C Unsupervised domain adaptation baselines
Importantly, following [42] we enforce all domain discriminators to be 1-Lipschitz. This allows us to approximate the empirical Wasserstein distance as:
where are the estimated Wasserstein distances of each of the anchor embeddings of a given level .
@@
SSTDA [7]: In the original work, SSTDA [7] proposes to couple what they called Local SSTDA with Global SSTDA. Local SSTDA refers to the standard frame-level alignment loss. To adapt this to our multi-resolution setup, we follow the same strategy as in our adaptation of DANN, applying adversarially training a classifier that learns to discriminate the domain that each of the frames/embeddings comes from.
@@
Global SSTDA, in contrast, refers to a sequential domain prediction loss. This splits the original source and target domain into 2 clips of sequential frames. They then apply a domain attentive temporal pooling [7] to first weight the frames according to their entropy, and then pool them into a single compact clip embedding. The source and target clip embeddings are then randomly shuffled, and finally passed through a domain classifier that learns to predict the permutation that these clips form (see the original work [7] for more details). To adapt this methodology to our multi-resolution setup, we apply this idea for every level independently. Concretely, we split into two segments the domain-invariant feature representations of domains and for a given level , yielding and . Then, following the original proposal, we apply domain attentive temporal pooling to compute segment-level representations and . Finally, we apply the sequential domain prediction on these segment embeddings of a given resolution level . For this, we shuffle them randomly into – e.g., – and train a level-wise domain classifier to predict which of the possible permutations was used.
@@
TranSVAE [65]: This work is one of the latest state-of-the-art methods on domain adaptation for video action recognition. This is especially interesting given the innovative disentanglement-based approach. This approach, concretely, leverages the use of a Variational Autoencoder that learns two types of representations: 1) Frame (clip) level representations with dynamic information 2) Static video level representation that includes domain-specific information. The main goal of this model is to enforce domain invariance on the frame-level representations. For this, they leverage several domain adversarial losses (see the original [65] for more details). Finally, it leverages a classification head to aggregate the domain invariant frame-level representations to perform a single video-level prediction.
@@
Adapting this work to our target task, Temporal Action Localization, is however nontrivial. Notice that we deal with sparse detection setups, so it is not enough to transform the goal of TranSVAE into a frame-level classification problem. We find this Action Segmentation to be ineffective for our chosen benchmarks. Consequently, we propose instead to substitute their original classification head for the SGP pyramid and classification/localization head that we describe in Sec. III.C and Sec. III.D. Hence, we utilize the domain invariant representations computed by TranSVAE as input of the multi-resolution pyramid and later prediction head. This makes this model architecture adaptable to our proposed experimental setup, and comparable to SADA.
Appendix D Extended results and ablations
@@
In this section, we extend the ablation studies presented in Sec. IV.C of the main text by analyzing two additional scenarios. Concretely, we compare the performance of SADA with the adaptations of different UDA methods, and with different variations of our loss, when evaluated in scenarios S2 and S4.
D.1 Domain adaptation methods
@@
In Table S5 we present the comparison of our proposed method to the other considered UDA baselines. Concretely, the first scenario (left) refers to the white new videos, while the second scenario (right) refers to the white old videos. Similarly to Sec. IV.C of the main text, we observe that SADA improves the results of all the considered baselines for both scenarios. More specifically, SADA attains a performance boost of up to mAP in the first scenario and in the second. Moreover, our method improves by mAP and mAP the performance of the second best-performing model in both scenarios, respectively. Interestingly, in both cases, the second best-performing method is DANN[1] which considerably improves other more complex methods like SSTDA[7]. Another interesting note is that WDGRL[42] presents a considerable variability in the performance. While this is the third best-performing method in the first scenario, its performance drops drastically in the second attaining only a mAP. This contrasts with the performance of similar methods like ADDA[40] or DANN[1].
[…]

### v3

@@ B.2 CharadesEgo scenario
Third-to-egocentric setup: Based on this dataset, we are the first to present a benchmark on domain adaptation for Temporal Action Localization. Our presented setup, concretely, explores the adaptation of third-person videos to egocentric ones. This sometimes involves an extreme domain shift as depicted in Fig. D. In fact we argue this domain shift can be so extreme that would not be suitable for image-based model benchmarking given that one of the views might lack necessary information that is only contained in the other view. For instance, a person awekens in bed could be indiscernible from a person looking at the ceiling if only one single frame is considered. Nevertheless, our proposed benchmark deals with video-inputs, which should provide sufficient context from the other frames to infer the true action, making this challenging setup more feasible.
To be more precise on our proposed benchmarking setup, we follow the work of [64] and define the third-person videos as Source domain and the egocentric ones as Target domain. This setup presents an important limitation that we find is critical to make this benchmark viable in this complex task. Concretely, observe that class actions are mostly defined as NLP descriptions, which creates very concrete actions, making it cumbersome to adapt classes across domains. We argue that one of the main reasons is the lack of sufficient data for each of the actions. Consequently, we rely on the originally extended annotations from [50] and propose to use instead the verbs of each of the corresponding action labels. This is, for a given action holding some clothes, we consider this action with its verb hold. One final consideration is that this still results in underrepresented verb classes which we fix following a similar approach to that presented by [57] and consider only the 10 majority classes being take, put, sit, walk, hold, stand, open, drink, smile and laugh. This still results in a very challenging setup as observed in Tab. 2, which indicates that none of the state-of-the-art methods tested attain more than mAP. Moreover, this setup retains of the original annotations.
C Unsupervised domain adaptation baselines
In this section, we describe in greater detail the different baselines that we use in Sec. 4.3 to establish a fair comparison with our work, SADA. Concretely, we first describe the adaptations to our architecture of multiple UDA methods like DANN [18], ADDA [55], WDGRL [47], MSTN [59], FGDA [20] or DRDA [24]. We then describe the adaptation of the closest related work to SADA, SSTDA [8] and for comparison purposes, the adaptation of one of the latest video action recognition methods, TranSVAE [58] to the task of Temporal Action Localization.
@@ C Unsupervised domain adaptation baselines
Importantly, following [47] we enforce all domain discriminators to be 1-Lipschitz. This allows us to approximate the empirical Wasserstein distance between the source and target domain as:
where for each of the anchor embeddings and resolution level .
@@
FGDA [20]: In the original work of [20], the authors argue about the equilibrium issues of existing adversarial methods that attempt to reduce the divergence between feature distributions. In this context, they propose a Feature Gradient Distribution Alignment (FGDA) approach that attempts to align the resulting gradients across the Source and Target domain, using a simple Gradient discriminator module. In our setup, this adaptation is relatively straightforward, and implies only minimal changes with respect to our previously described DANN [18] baseline. Concretely, in DANN we fed the intermediate anchor representations and to the domain discriminator . In this case, we leverage instead their respective gradients and . These are fed as input to a gradient domain discriminator that has the same architecture as that used in DANN, which aims to discern the domain that produced these gradients. This results in the following loss:
Note that this loss is trained in a standard adversarial fashion by incorporating a GRL module that inverts the gradients of the discriminator.
DRDA [24]: This work observes that the features of different class labels often expand outwards in multiple directions, forming an overall radial pattern. They argue that transferring this discriminative pattern across domains can be beneficial for performance. To adapt this method to our proposed architecture, we follow a similar approach to that described in the original work and implement their Global Transformation and a Local Refinement. More specifically, the Global Transformation performs a similar process to that described for the MSTN[59], but with more global, level-wise, centroids. Hence, we compute the set of global centroids and , which are defined as the mean feature embedding for each of the given resolution levels. Then, in order to reduce the expected mean feature shift, we use an MSE loss to reduce to distance.
For the Local Refinement, we essentially follow the same formulation described in detail in the original work [24]. The only adaptation that we make is that the original work operates on a single-resolution level. Thus, for our baseline we define the class-level wise centroids as described in Eq. 16 - 19. Then, we follow exactly their formulation to define the radial structure loss described in [24] using our computed local centroids and . Importantly, we define the local refinement loss for each of the levels independently and then average across levels to compute the final refinement loss.
@@
SSTDA [8]: In the original work, SSTDA [8] proposes to couple what they called Local SSTDA with Global SSTDA. Local SSTDA refers to the standard frame-level alignment loss. To adapt this to our multi-resolution setup, we follow the same strategy as in our adaptation of DANN, applying adversarially training a classifier that learns to discriminate the domain that each of the frames/embeddings comes from. Global SSTDA, in contrast, refers to a sequential domain prediction loss. This splits the original source and target domain into 2 clips of sequential frames. They then apply a domain attentive temporal pooling [8] to first weight the frames according to their entropy, and then pool them into a single compact clip embedding. The source and target clip embeddings are then randomly shuffled, and finally passed through a domain classifier that learns to predict the permutation that these clips form (see the original work [8] for more details). To adapt this methodology to our multi-resolution setup, we apply this idea for every level independently. Concretely, we split into two segments the domain-invariant feature representations of domains and for a given level , yielding and . Then, following the original proposal, we apply domain attentive temporal pooling to compute segment-level representations and . Finally, we apply the sequential domain prediction on these segment embeddings of a given resolution level . For this, we shuffle them randomly into – e.g., – and train a level-wise domain classifier to predict which of the possible permutations was used.
@@
TranSVAE [58]: This work is one of the latest state-of-the-art methods on domain adaptation for video action recognition. This is especially interesting given the innovative disentanglement-based approach. This approach, concretely, leverages the use of a Variational Autoencoder that learns two types of representations: 1) Frame (clip) level representations with dynamic information and 2) Static video level representation that includes domain-specific information. The main goal of this model is to enforce domain invariance on the frame-level representations. For this, they leverage several domain adversarial losses (see the original [58] for more details). Finally, it leverages a classification head to aggregate the domain invariant frame-level representations to perform a single video-level prediction.
@@
Adapting this work to our target task, Temporal Action Localization, is however nontrivial. Notice that we deal with sparse detection setups, so it is not enough to transform the goal of TranSVAE into a frame-level classification problem. We find this Action Segmentation to be ineffective for our chosen benchmarks. Consequently, we propose instead to substitute their original classification head for the SGP pyramid and classification/localization head that we describe in Sec. 3.3 and 3.4. Hence, we utilize the domain invariant representations computed by TranSVAE as input of the multi-resolution pyramid and later prediction head. This makes this model architecture adaptable to our proposed experimental setup, and comparable to SADA.
D Extended results and ablations
@@
In this section, we expand the ablation studies presented in Sec. 4.3 by analyzing three additional scenarios. Concretely, we compare the performance of SADA with the adaptations of different UDA methods, and with different variations of our loss, in S1, S2 and S4.
D.1 Extened comparison with existing UDA methods
@@
In Tab. 3 of the main text we compared the performance of our proposed method to existing UDA baselines when evaluated on scenario S3. To gain further insights on this study, in Tab. E we extend this analysis to 3 further scenarios –i.e., scenarios S1, S2 and S4. In this regard, similarly to Sec. 4.3, we observe that SADA improves the results of all the considered baselines for all these scenarios. More specifically, SADA attains a performance boost of up to mAP, mAP, and , on the first, second and third considered scenario, respectievely. Moreover, our method improves by mAP, mAP and mAP the performance of the second best-performing model in these scenarios, respectively. Interestingly, in most of these cases, the second best-performing method is DANN[19] which considerably improves other more complex methods like SSTDA[8]. Another interesting note is that WDGRL[47] and DRDA[24] present a considerable variability in the performance. For instance, while WDGRL is the third best-performing method in the second scenario, its performance drops drastically in the first and third scenarios attaining only a mAP and mAP, respectively. Similarly, FGDA[24] nearly achieves state-of-the-art in S1 but considerably drops in S4. This contrasts with the performance of similar methods like ADDA[55] or DANN[19] that show a much more robust performance across the different scenarios.
S1 | ||||
@@
DANN [19] | 30.87 | 27.45 | 18.29 | 25.97 |
ADDA [55] | 30.55 | 27.09 | 18.17 | 25.27 |
WDGRL [47] | 24.85 | 21.87 | 13.68 | 20.13 |
FGDA [20] | 29.34 | 26.99 | 17.54 | 24.62 |
DRDA [24] | 31.50 | 27.97 | 18.79 | 26.08 |
MSTN [59] | 30.88 | 27.44 | 18.25 | 25.52 |
SSTDA [8] | 29.92 | 26.49 | 17.55 | 24.65 |
TranSVAE [58] | 25.88 | 21.74 | 13.12 | 20.25 |
Ours (SADA) | 31.60 | 28.22 | 18.98 | 26.27 |
S2 | ||||
Model | mAP {10,30,50}% | Avg | ||
DANN [19] | 31.32 | 28.55 | 20.10 | 26.65 |
ADDA [55] | 30.16 | 27.58 | 19.72 | 25.82 |
WDGRL [47] | 24.77 | 21.96 | 14.81 | 20.51 |
FGDA [20] | 25.20 | 22.72 | 15.05 | 20.99 |
DRDA [24] | 28.78 | 26.13 | 17.58 | 24.16 |
MSTN [59] | 30.17 | 27.32 | 18.55 | 25.34 |
SSTDA [8] | 29.58 | 26.81 | 18.36 | 24.92 |
TranSVAE [58] | 28.31 | 25.67 | 16.72 | 23.57 |
Ours (SADA) | 31.54 | 28.77 | 20.22 | 26.84 |
S4 | ||||
Model | mAP {10,30,50}% | Avg | ||
DANN [19] | 34.11 | 30.48 | 20.75 | 28.44 |
@@
D.2 Extended ablation over variants of the SADA loss
S1 | ||||||
Local | Global | Bkg | mAP {10,30,50}% | Avg | ||
30.74 | 27.23 | 18.07 | 25.34 | |||
30.87 | 27.45 | 18.29 | 25.97 | |||
30.97 | 27.64 | 18.39 | 25.67 | |||
30.36 | 26.82 | 18.01 | 25.06 | |||
30.82 | 27.57 | 18.30 | 25.56 | |||
30.30 | 26.98 | 17.85 | 25.04 | |||
31.60 | 28.22 | 18.98 | 26.27 | |||
S2 | ||||||
Local | Global | Bkg | mAP {10,30,50}% | Avg | ||
29.65 | 26.86 | 19.14 | 25.21 | |||
31.32 | 28.55 | 20.10 | 26.65 | |||
29.94 | 27.39 | 19.40 | 25.57 | |||
30.86 | 28.02 | 19.84 | 26.57 | |||
30.84 | 27.98 | 19.37 | 26.06 | |||
29.77 | 27.10 | 18.70 | 25.19 | |||
31.54 | 28.77 | 20.22 | 26.84 | |||
S4 | ||||||
Local | Global | Bkg | mAP {10,30,50}% | Avg | ||
34.41 | 30.58 | 21.00 | 28.66 | |||
34.11 | 30.48 | 20.75 | 28.44 | |||
34.46 | 30.85 | 20.90 | 28.73 | |||
34.63 | 31.08 | 21.26 | 28.99 | |||
33.47 | 29.91 | 20.17 | 27.85 | |||
32.84 | 29.06 | 19.41 | 27.10 | |||
34.86 | 31.16 | 21.46 | 29.16 |
Closely related to Sec. 4.3 of the main text, in Tab. F we extend the ablation over different variants of our loss on scenarios S1, S2 and S4. Observe in this table that our main insights from Sec. 4.3 of the main text are mostly confirmed. Concretely, one of our main observations was that aligning the complete global distribution was generally preferable over aligning the partial background distribution only. This observation is satisfied in most of the scenarios that we study –with the exception of scenario S2. Similarly, in Sec. 4.3 of the main text we argued that combining alignment loss that presented overlaps yielded also a decrease in performance. More in detail, this happens when we couple either the local or the background loss with a global alignment approach. In these cases, the embeddings have a double (simultaneous) alignment objective, which we argue is not desirable. As observed in Tab. F, this observation is also mostly satisfied in these scenarios as the combination of these losses yields a performance degradation in 7 of the 8 cases studied in this paper.
E Ablation of the effect of the hyper parameters
As described in Eq. 9, we define our SADA loss using a set of hyperparameters , where each of the parameters controls the influence of a given resolution level in the overall loss. Given the relevance of these parameters, below we ablate over different variations to showcase the effect that they have in the final performance of the model. Concretely, we define the three following scenarios:
1.
All levs 1: This scenario sets all the parameters , keeping the contribution of each of the levels the same.
2.
First 3 levs 1: This scenario sets the first half of the levels to 1 and the rest to 0. This is and .
3.
Last 3 levs 1: This scenario sets the first half of the levels to 0 and the rest to 1. This is and .
All in all, this study attempts to clarify how sensitive is our model to this design decision. In this regard, these results (see Tab. G) indicate a mild sensitivity to this choice. Specifically, we observe that the second best-performing method attains an absolute difference over the optimal choice of only mAP and mAP for the two considered scenarios S1 and S3, respectively. Moreover, we observe that the worst performing naive strategy still outperforms in both scenarios 4 out of the 6 tested UDA baselines (see Tab. 3).
S1 | ||||
Strategy | mAP {10,30,50}% | Avg | ||
All levs 1 | 30.60 | 27.36 | 18.41 | 25.45 |
First 3 levs 1 | 31.05 | 27.64 | 18.30 | 25.66 |
Last 3 levs 1 | 30.95 | 27.59 | 18.56 | 25.70 |
"Optimal choice" | 31.60 | 28.22 | 18.98 | 26.27 |
S3 | ||||
Strategy | mAP {10,30,50}% | Avg | ||
All levs 1 | 32.42 | 29.59 | 19.34 | 27.11 |
First 3 levs 1 | 30.96 | 27.48 | 18.29 | 25.58 |
Last 3 levs 1 | 30.54 | 27.19 | 18.28 | 25.34 |
"Optimal choice" | 32.69 | 29.17 | 19.72 | 27.19 |
F Study of the class embedding
S1 | ||||||
Strategy | mAP {10,20,30,40,50}% | Avg | ||||
One-hot | 31.31 | 29.97 | 27.84 | 24.28 | 18.81 | 26.44 |
Random emb | 31.38 | 30.01 | 27.94 | 24.13 | 18.70 | 26.43 |
Sinusoidal [56] | 31.10 | 29.78 | 27.71 | 23.99 | 18.56 | 26.23 |
Learnable | 31.68 | 30.32 | 28.37 | 24.55 | 19.09 | 26.80 |
S3 | ||||||
Strategy | mAP {10,20,30,40,50}% | Avg | ||||
One-hot | 32.29 | 31.09 | 28.58 | 24.88 | 19.56 | 27.28 |
Random emb | 30.71 | 29.50 | 27.43 | 23.78 | 18.52 | 25.99 |
Sinusoidal [56] | 29.93 | 28.57 | 26.44 | 22.87 | 17.01 | 24.96 |
Learnable | 32.69 | 31.49 | 29.17 | 25.51 | 19.72 | 27.72 |
One of the main contributions of our work is to adversarially align distributions in a class-wise fashion. Intuitively, this requires that our level-wise domain discriminator knows the class distribution that it is aligning. In this regard, in Sec. 3.5 we propose to concatenate to every anchor a learnable embeddings of its corresponding class (see Eq. 7-8). In this ablation, we empirically justify our choice. For this, we compare our approach against several non-learnable alternatives to encode a given class : a naive one-hot encoding, a random class-wise dense initialization, and the sinusoidal encoding originally proposed by [56].
In Tab. H we show the experimental results of each of the variants tested in S1 and S3. These indicate that a naive one-hot encoding of a class is the best-performing non-learnable strategy, obtaining considerable improvements over the other two tested non-learnable baselines. This improvement is especially prominent in the first scenario, boosting the performance over the other non-learnable strategies by mAP and mAP, respectively. Moreover, to our surprise, the popular Sinusoidal encoding [56] proves to be the worst-performing method, being consistently outperformed by even the random dense encoding. Finally, we highlight that in both scenarios the use of a learnable class embedding yields the best results, improving by mAP and mAP the performance of the one-hot encoding strategy, which justifies our choice.
G Ablation per class
take | put | wash | open | close | insert | turn-on | cut | turn-off | pour | |
take | put | wash | open | close | insert | turn-on | cut | turn-off | pour | |
In this section, we complement the analysis from Sec. 4.2 with the detailed class-wise metrics. Concretely, in Tab. I and Tab. J we show the respective class-wise mAP scores of the 10 considered classes on both S1 and S3. For the analysis, we also include the results obtained with the source-only variant of our model as well as our chosen DANN[18] baseline.
Observe that in S3 (see Tab. I) our model attains the best class-wise performance in 5 of the classes, while DANN [18] does so on 3, and the source-only model on 2. In contrast, in S1 (see Tab. J) our method obtains a much clearer improvement over the chosen baselines, yielding the best results in 8 of the 10 classes. Overall, we can observe that our method performs very well in the 3 majority classes of both scenarios, where we highlight the absolute improvement of SADA over DANN [18] of mAP for the class put in S3. Moreover, our method fails to improve the performance of action pour in both scenarios, which we attribute to the lack of sufficient data to operate on our proposed methodology. We also highlight that as observed in Sec. 4.3, S1 presents a more challenging setup, degrading the performance in other actions such as open, close, insert or cut. This might indicate the existence of intrinsic qualitative aspects that harden the adaptation when dealing with old videos. This is not the case in Scenario 3, as aside from the aforementioned pour segments, it only fails to achieve the best results for the turn-on action.
H Ablation of the effect of background anchors
In this work we focus on applying domain adaptation over anchor-based localization methods, the current state-of-the-art architectures on datasets like EpicKitchens100 [13]. The success of these architectures is normally explained by the generation of a very considerable number of anchors, largely surpassing the number of actual ground-truth labels to predict. As described in Sec. 3, these are the excess anchors –i.e., background anchors– that are not directly matched with a GT segment, consequently training the model to predict class 0 instead. In other words, in these cases we hope for the model to learn to identify these anchor embeddings as background, not predicting any action class during inference. This, however, is a cumbersome issue that most of the anchor-based methods face, and which results in a challenging model transfer to a new data domain. It is not clear, however, weather the use domain adaptation methods effectively help to mitigate the confusion that these embeddings induce, which we largely attribute to a high intra-class variance and a low inter-class variance with the other action classes. Hence, even though we leave a more thorough analysis as future work, in this ablation we aim to provide minimal proof of how our method partially mitigates this issue, and also of the potential that these methods would have could we manage to solve this issue.
H.1 Masking out the anchors during inference
For this, in this section we firstly extend to scenario S1 the experiment presented in Sec. 4.3 of the main text (see Tab. 5), where we propose an experiment where mask out the set of all the background anchors during inference. Put differently, we leave all the training routines unchanged – thus considering all the anchors– while not considering the background anchors at all during inference. This measures the predicting potential of a model that is able to perfectly deal with these embeddings. Observe in Tab. K that in this scenario, our proposed method SADA has a potential increase of mAP if we mask all the background anchors during inference. This contrasts with the mAP increase in the source-only version, which shows that SADA reduces this gap by mAP. In other words, SADA partially improves the transferability of these anchors across domains.
Method | Mask bkg anchors | mAP {10,30,50}% | Avg | Perf. gap | ||
Ours(src-only) | 30.74 | 27.23 | 18.07 | 25.38 | - | |